1. Consider an air-filled parallel plate capacitor with plate area A = 37.4 cm2 and gap width d = 0.7 mm.

A. What is its capacitance?

B. The capacitor was connected to a 40-V battery, and disconnected after charging. What is the charge Q drawn from the battery?

2. The gap of the capacitor described in Problem #1 is now filled by the material with the dielectric constant of 4.2.

A. What is its capacitance?

B. The capacitor was connected to a 40-V battery, and disconnected after charging. What is the charge Q drawn from the battery?
Answer #1

Solution 1)

Part A) We know, capacitance C= AEo/d

C= (37.4*10^-4)m^2 *(8.85*10^-12)/(0.7*10^-3)

C= 0.0469 F (Ans)

======

Part b)

Here, V=40 V

We know, Q= CV= 0.0469*40= 1.878 C (Ans)

======

Acc to HOMEWORKLIB RULES, I can answer only one at a time. Good luck!:)